OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). a = 3b - 4 a + b =16

OpenStudy (anonymous):

u should try this

OpenStudy (anonymous):

substract 2 of 1 (2) -(1) a + b -a =16 -3b+4 b+3b=20 b=20/4 b=5 solve (2) or (1) with b a+5=16 a=11

OpenStudy (anonymous):

3b-4+b=16 4b=20 b=5 a+b=16 a+5=16 a=11 (11,5)